A 2.0L sample of gas at 27C and 800 torr is cooled to -123C. If the pressure remains constant, what is its new volume?
1
Expert's answer
2012-12-10T10:49:07-0500
if pressure p=const, then
V/T is also constant

V1/T1 = V2/T2

V2 = V1 * T2/T1 = 2 L * (273-123)K/(273+27)K = 2L * 150K/300k = 2L/2 = 1L
new volume is 1 L.